# Application of Analytical Hierarchical Process and its Variants on Remote Sensing Datasets

### v1

@@ 2.3.1. Nested AHP
Using a pairwise comparison matrix, we assessed the relationship between individual parameter values. Table 2 shows the pairwise comparison of Population Density and the weight each category has, e.g. PD of ¿4036 people per km2 has a 9 times greater effect on vulnerability than when PD was ¡865 people per km2. The pairwise matrices for the other factors alongside the explanations of the comparison indices were provided in the Supplemental Information section.
Table 2. Nested AHP: Pairwise Comparison for Individual Parameter Values for Population Density
Population Density / km2 ¿4036 2680-4036 1600-2680 865-1600 ¡865
¿4036 1 4 6 8 9
@@
Table 3. Nested AHP: Priority Vector of Subweights for each factor
@@
    (9) $W=\begin{bmatrix}w_{1}&w_{2}&w_{3}&w_{4}&w_{5}&w_{6}\end{bmatrix}$
@@
    (10) $X(p)=\begin{bmatrix}x_{1}(p)&x_{2}(p)&x_{3}(p)&x_{4}(p)&x_{5}(p)&x_{6}(p)\end{bmatrix}$
@@
    (11) $S=\begin{bmatrix}s_{1}(1)&s_{2}(1)&s_{3}(1)&s_{4}(1)&s_{5}(1)&s_{6}(1)\\ s_{1}(2)&s_{2}(2)&s_{3}(2)&s_{4}(2)&s_{5}(2)&s_{6}(2)\\ s_{1}(3)&s_{2}(3)&s_{3}(3)&s_{4}(3)&s_{5}(3)&s_{6}(3)\\ s_{1}(4)&s_{2}(4)&s_{3}(4)&s_{4}(4)&s_{5}(4)&s_{6}(4)\\ s_{1}(5)&s_{2}(5)&s_{3}(5)&s_{4}(5)&s_{5}(5)&s_{6}(5)\end{bmatrix}$
@@
(12) $P=\begin{bmatrix}0&1&0&0&0&0\\ 0&0&0&0&1&0\\ 1&0&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&1\end{bmatrix}$
@@
(13) $S\cdot P=\begin{bmatrix}s_{1}(x_{1}(p))&s_{2}(x_{2}(p))&...&s_{6}(x_{6}(p))\end{bmatrix}$
This operation selects the subweights corresponding to the actual values of the parameters at pixel p. Finally, we multiplied this resulting subweight vector element-wise with the AHP weight vector W and summed the results to obtain the predicted vulnerability for each pixel using Nested AHP V(p):
(14) $V(p)=W\cdot(S\cdot P)^{T}$
@@
2.3.2. Fuzzy AHP
Fuzzy Analytic Hierarchy Process (Fuzzy AHP), as introduced by Saaty (1982)(21, ) is an extension of the classic AHP that incorporates the fuzziness (probabilistic alterations/error) associated with human judgment. By using fuzzy numbers in the rankings, Fuzzy AHP accommodates these uncertainties and thus provides more reliable results.
@@ 2.3.2. Fuzzy AHP
2.3.3. ANP
One of the main limitations of AHP is its inability to adequately capture the interdependencies among criteria and alternatives. To address this issue, Saaty (2006)(22, ) developed ANP, which extends the AHP methodology by accommodating feedback loops and interdependencies among criteria and alternatives. Unlike AHP, which assumes a hierarchical structure, ANP allows for the representation of complex relationships and feedback loops in decision networks. The method utilizes a supermatrix to integrate both local and global influences, providing a more realistic representation of environments. ANP is widely used to represent interdependent relationships, making it suitable for decision problems with interconnected elements. ANP introduces the concept of contextual weights, enabling decision makers to adjust the importance of criteria and alternatives based on the context of the decision problem. This flexibility enhances the adaptability of ANP to various situations.
In a complex system such as the geophysical environment, it is expected that AHP factors will be correlated to one another. While AHP assumes strict independence between criteria at different levels of the hierarchy, ANP allows for both dependence and independence relationships (Table 4) as described previously by Poh and Liang (2017)(18, ); this flexibility better captures the intricacies of decision networks across natural systems.
Table 4. Vulnerability ANP Supermatrixa
@@ 2.3.3. ANP
  • Inner dependence (As defined by Poh and Liang (2017)(18, )) matrix of criteria, $W_{22}$;
@@
We used the Python library pyanp created by Adams, et al. (2020)(1, ) to run the ANP Algorithm on the Supermatrix.
2.3.4. 1-N AHP
@@ 2.3.5. Fuzzy 1-N AHP
3. Results
3.1. River Ganga vulnerability based on AHP
Given our pairwise matrix, we obtained the importance weights of parameters as 0.408, 0.268, 0.138, 0.100, 0.058, 0.028 for Population Density (PD), LULC, Rainfall (RAIN), Drainage Density (DD), Slope, Land Surface Temperature (LST) respectively. We multiplied the pixel values for each parameter with their weights to get our final output layer, which represented the vulnerability map of the River Ganga (Figure 3).
Figure 3. AHP: Weighted Overlay Vulnerability Map & Histogram
Figure 3. These figures show the AHP’s vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Figure 3a. shows the exact locations and values, while figure 3b. shows the count for each bin range of the vulnerability in histogram form.
AHP vulnerability scores ranged from a minimum of 1.001 to a maximum of 4.470 with 83.7% of the area being categorized as having extremely low or low vulnerability and 3.5% of the area is highly or extremely high vulnerability (Figure 3b). Tier 2 cities (i.e., cities with a population between 50,000 and 99,999 people) on the banks of river Ganga such as Varanasi, Prayagraj/Allahabad, and Kanpur were clearly highlighted with high and extremely high AHP vulnerability scores (Figure 3a). Indeed, urban settings that had population densities $\geq$ 1,100 people km2 typically had vulnerability scores $\geq$ 2.736. Alternatively, the areas with vulnerability scores $\leq$ 1.694 were characterized by forest or cropland regions with low human population densities. The annual rainfall in the high elevation regions of the river Ganga generally increased the vulnerability of those regions.
3.2. Nested AHP and ANP
@@
The distribution of vulnerability scores produced with both Nested AHP and ANP were markedly different from the distribution produced with AHP (Figure 4 and Figure 5). Specifically, Nested AHP resulted in 96.7% of the area scored as extremely low and low vulnerability, which was skewed with a right-tail range that extended 1.8-fold further than AHP. Similarly, ANP exhibited a right-tailed distribution with low vulnerability scores. Despite that, the Tier 2 cities had high vulnerability scores for both ANP and Nested AHP, except for Kanpur, which saw a major reduction in the vulnerability scores. These differences were due to the nonlinear response of factor weights to vulnerability rankings in both methods (Figure 6). Thus, these methods resulted in overall focused, yet dampened, vulnerability-score maps compared to AHP. Despite the additional insights these analyses revealed, the fundamental issue of the extent to which the defined AHP factors adequately encompass the vulnerability space remained.
@@
Figure 4. Nested AHP: Ganga Vulnerability Map & Histogram
Figure 4. These figures show the Nested AHP’s vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Figure 4a. shows the exact locations and values, while figure 4b. shows the count for each bin range of the vulnerability in histogram form.
Figure 5. ANP: Ganga Vulnerability Map & Histogram
Figure 5. These figures show the ANP’s vulnerability score for the scope of our analysis, the lower values are scored as the least vulnerable areas while the ones with higher scores were ranked as the most vulnerable. Figure 5a. shows the exact locations and values, while figure 5b. shows the count for each bin range of the vulnerability in histogram form
Figure 6. Nested AHP, ANP & AHP Comparison
Figure 6. This graph shows the normalized amount each factor class 1, 2, 3, 4, 5 affects the overall weighted overlay for AHP, Nested AHP and ANP, where blue (circles) represents AHP, red (squares) represents Nested AHP and green (triangles) represents ANP.
3.3. 1-N AHP
When we ran 1-N AHP for the worst case, N = 0.33, we attributed more weightage to the unknown acute and chronic factors than for the average case, N = 0.165. However, due to the low probability of the presence of an acute or chronic factor, the vulnerability scores were lower for the worst case than the average case (Figure 7 and Figure 8). Even when we considered the distribution of the result values, we saw that the results for the worst case were more right skewed than the average case, which was more right skewed than the original AHP. From this we can gather that given our assumptions, there was no negative impact of considering unknown factors in our AHP analysis. However, 1-N would positively be able to capture the robustness of AHP results in other contexts, or if the assumptions were altered, and since our acute and chronic layers were purely random across the Ganga buffer, the change in vulnerability values were correlated to the values of the initial AHP results, and not to any specific remote sensing or other GIS data layers in the analysis.
Figure 7. 1 - N AHP Worst case (N = 0.33): Ganga Vulnerability Map & Histogram
Figure 7. These figures show the 1 - N AHP worst case’s vulnerability score when we consider the worst case (N=0.33); the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Figure 7a. shows the exact locations and values, while figure 7b. shows the count for each bin range of the vulnerability in histogram form
Figure 8. 1 - N AHP Average case (N = 0.165): Ganga Vulnerability Map & Histogram
Figure 8. These figures show the 1 - N AHP average case’s vulnerability score when we consider the average/expected case (N=0.165); the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Figure 8a. shows the exact locations and values, while figure 8b. shows the count for each bin range of the vulnerability in histogram form
3.4. Fuzzy AHP and Fuzzy 1-N AHP
@@
Table 5. Count of Rank Reversals for Fuzzy AHP out of 100,000 simulations
@@
The simulations with rank reversals were tested to determine their potential impact on vulnerability scores. This was done by running AHP on a sample of 500 from the 4,433 cases where a rank reversal occurred, and averaging all the output layers (Figure 9). Of these, areas that experienced changes in vulnerability due to rank reversals had high Drainage Density, or were areas that tended to receive large amounts of precipitation.
Figure 9. Fuzzy AHP Average: Ganga Vulnerability Map & Histogram
Figure 9. These figures show the Fuzzy AHP Average case’s vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Figure 9a. shows the exact locations and values, while figure 9b. shows the count for each bin range of the vulnerability in histogram form.
@@
For both the cases, we saw the distribution skewed to the left as compared to the original AHP (Figure 10 and Figure 11), and the vulnerability increased, which was more prominent for Case 2 (Figure 11). Due to the high weightage of the parameter, there was a increase in vulnerability scores in the areas with high population density.
Figure 10. Fuzzy AHP Case 1: Ganga Vulnerability Map & Histogram
Figure 10. These figures show the Fuzzy AHP Case 1 vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Figure 10a. shows the exact locations and values, while figure 10b. shows the count for each bin range of the vulnerability in histogram form
Figure 11. Fuzzy AHP Case 2: Ganga Vulnerability Map & Histogram
Figure 11. These figures show the Fuzzy AHP Case 2’s vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Figure 11a. shows the exact locations and values, while figure 11b. shows the count for each bin range of the vulnerability in histogram form
@@
Table 6. Count of First-order Rank Reversals for Fuzzy 1-N AHP out of 100,000 simulations
Reversals between Parameters Range of Original AHP results Rank Reversals
@@
Table 7. Count of Second-order Rank Reversals for Fuzzy 1-N AHP out of 100,000 simulations
Reversals between Parameters Range of Original AHP results Rank Reversals
@@
3.5. Comparison of analysis scores
@@
To visualize and compare the difference of the AHP layer with other approaches, we created 3 different map types (Figure 12). We created Map Type 1 by stretching the colors between the overall min max of all the different approaches, to compare the approaches among themselves. For Map Type 2 we stretched the colors between the local minimum and maximum value of each approach, to make the outliers for each individual map stand out. And for Map Type 3, we stretched the colors between the local 5 percentile and 95 percentile for each approach, to prevent the outliers from washing-out the differences in the map.
Figure 12. Original AHP & AHP Variants Comparison Charts
Figure 13. Original AHP & AHP Variants Comparison Charts (continued)
Figure 12. These figures show the comparison charts for the different variants of AHP with the original AHP across the Ganga buffer. We use three map types to highlight the differences between the approaches.
3.5.1. Difference between Nested AHP and ANP

### v2

@@ 2.3.1 Nested AHP
Using a pairwise comparison matrix, we assessed the relationship between individual parameter values. Table 2 shows the pairwise comparison of Population Density and the weight each category has, e.g. PD of >4036 people per km2 has a 9 times greater effect on vulnerability than when PD was <865 people per km2. The pairwise matrices for the other factors alongside the explanations of the comparison indices were provided in the Supplemental Information section.
Table 2: Nested AHP: Pairwise Comparison for Individual Parameter Values for Population Density
Population Density / km2 >4036 2680-4036 1600-2680 865-1600 <865
>4036 1 4 6 8 9
@@
Table 3: Nested AHP: Priority Vector of Subweights for each factor
@@
    $W=\begin{bmatrix}w_{1}&w_{2}&w_{3}&w_{4}&w_{5}&w_{6}\end{bmatrix}$ (9)
@@
    $X(p)=\begin{bmatrix}x_{1}(p)&x_{2}(p)&x_{3}(p)&x_{4}(p)&x_{5}(p)&x_{6}(p)\end{bmatrix}$ (10)
@@
    $S=\begin{bmatrix}s_{1}(1)&s_{2}(1)&s_{3}(1)&s_{4}(1)&s_{5}(1)&s_{6}(1)\\ s_{1}(2)&s_{2}(2)&s_{3}(2)&s_{4}(2)&s_{5}(2)&s_{6}(2)\\ s_{1}(3)&s_{2}(3)&s_{3}(3)&s_{4}(3)&s_{5}(3)&s_{6}(3)\\ s_{1}(4)&s_{2}(4)&s_{3}(4)&s_{4}(4)&s_{5}(4)&s_{6}(4)\\ s_{1}(5)&s_{2}(5)&s_{3}(5)&s_{4}(5)&s_{5}(5)&s_{6}(5)\end{bmatrix}$ (11)
@@
$P=\begin{bmatrix}0&1&0&0&0&0\\ 0&0&0&0&1&0\\ 1&0&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&1\end{bmatrix}$ (12)
@@
$S\cdot P=\begin{bmatrix}s_{1}(x_{1}(p))&s_{2}(x_{2}(p))&...&s_{6}(x_{6}(p))\end{bmatrix}$ (13)
This operation selects the subweights corresponding to the actual values of the parameters at pixel p. Finally, we performed the dot product of the AHP weight vector $W$ with the transpose of this resulting subweight vector $S.P$ to obtain the predicted vulnerability for each pixel using Nested AHP $V(p)$:
$V(p)=W\cdot(S\cdot P)^{T}$ (14)
@@
2.3.2 Fuzzy AHP
Fuzzy Analytic Hierarchy Process (Fuzzy AHP), as introduced by Saaty (1982) is an extension of the classic AHP that incorporates the fuzziness (probabilistic alterations/error) associated with human judgment. By using fuzzy numbers in the rankings, Fuzzy AHP accommodates these uncertainties and thus provides more reliable results.
@@ 2.3.2 Fuzzy AHP
2.3.3 ANP
One of the main limitations of AHP is its inability to adequately capture the interdependencies among criteria and alternatives. To address this issue, Saaty (2006) developed ANP, which extends the AHP methodology by accommodating feedback loops and interdependencies among criteria and alternatives. Unlike AHP, which assumes a hierarchical structure, ANP allows for the representation of complex relationships and feedback loops in decision networks. The method utilizes a supermatrix to integrate both local and global influences, providing a more realistic representation of environments. ANP is widely used to represent interdependent relationships, making it suitable for decision problems with interconnected elements. ANP introduces the concept of contextual weights, enabling decision makers to adjust the importance of criteria and alternatives based on the context of the decision problem. This flexibility enhances the adaptability of ANP to various situations.
In a complex system such as the geophysical environment, it is expected that AHP factors will be correlated to one another. While AHP assumes strict independence between criteria at different levels of the hierarchy, ANP allows for both dependence and independence relationships (Table 4) as described previously by Poh and Liang (2017); this flexibility better captures the intricacies of decision networks across natural systems.
Table 4: Vulnerability ANP Supermatrixa
@@ 2.3.3 ANP
  • Inner dependence (As defined by Poh and Liang (2017)) matrix of criteria, $W_{22}$;
@@
We used the Python library pyanp created by Adams et al. (2020) to run the ANP Algorithm on the Supermatrix.
2.3.4 1-N AHP
@@ 2.3.5 Fuzzy 1-N AHP
3 Results
3.1 River Ganga vulnerability based on AHP
Given our pairwise matrix, we obtained the importance weights of parameters as 0.408, 0.268, 0.138, 0.100, 0.058, 0.028 for Population Density (PD), LULC, Rainfall (RAIN), Drainage Density (DD), Slope, Land Surface Temperature (LST) respectively. We multiplied the pixel values for each parameter with their weights to get our final output layer, which represented the vulnerability map of the River Ganga (Fig 3).
Fig 3. These figures show the AHP’s vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
AHP vulnerability scores ranged from a minimum of 1.001 to a maximum of 4.470 with 83.7% of the area being categorized as having extremely low or low vulnerability and 3.5% of the area is highly or extremely high vulnerability (Fig 3b). Tier 2 cities (i.e., cities with a population between 50,000 and 99,999 people) on the banks of river Ganga such as Varanasi, Prayagraj/Allahabad, and Kanpur were clearly highlighted with high and extremely high AHP vulnerability scores (Fig 3a). Indeed, urban settings that had population densities $\geq$ 1,100 people km2 typically had vulnerability scores $\geq$ 2.736. Alternatively, the areas with vulnerability scores $\leq$ 1.694 were characterized by forest or cropland regions with low human population densities. The annual rainfall in the high elevation regions of the river Ganga generally increased the vulnerability of those regions.
3.2 Nested AHP and ANP
@@
The distribution of vulnerability scores produced with both Nested AHP and ANP were markedly different from the distribution produced with AHP (Fig 4 and Fig 5). Specifically, Nested AHP resulted in 96.7% of the area scored as extremely low and low vulnerability, which was skewed with a right-tail range that extended 1.8-fold further than AHP. Similarly, ANP exhibited a right-tailed distribution with low vulnerability scores. Despite that, the Tier 2 cities had high vulnerability scores for both ANP and Nested AHP, except for Kanpur, which saw a major reduction in the vulnerability scores. These differences were due to the nonlinear response of factor weights to vulnerability rankings in both methods (Fig 6). Thus, these methods resulted in overall focused, yet dampened, vulnerability-score maps compared to AHP. Despite the additional insights these analyses revealed, the fundamental issue of the extent to which the defined AHP factors adequately encompass the vulnerability space remained.
@@
Fig 4. These figures show the Nested AHP’s vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
Fig 5. These figures show the ANP’s vulnerability score for the scope of our analysis, the lower values are scored as the least vulnerable areas while the ones with higher scores were ranked as the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
Fig 6. This graph shows the normalized amount each factor class 1, 2, 3, 4, 5 affects the overall weighted overlay for AHP, Nested AHP and ANP, where blue (circles) represents AHP, red (squares) represents Nested AHP and green (triangles) represents ANP.
3.3 1-N AHP
When we ran 1-N AHP for the worst case, N = 0.33, we attributed more weightage to the unknown acute and chronic factors than for the average case, N = 0.165. However, due to the low probability of the presence of an acute or chronic factor, the vulnerability scores were lower for the worst case than the average case (Fig 7 and Fig 8). Even when we considered the distribution of the result values, we saw that the results for the worst case were more right skewed than the average case, which was more right skewed than the original AHP. From this we can gather that given our assumptions, there was no negative impact of considering unknown factors in our AHP analysis. However, 1-N would positively be able to capture the robustness of AHP results in other contexts, or if the assumptions were altered, and since our acute and chronic layers were purely random across the Ganga buffer, the change in vulnerability values were correlated to the values of the initial AHP results, and not to any specific remote sensing or other GIS data layers in the analysis.
Fig 7. These figures show the 1 - N AHP worst case’s vulnerability score when we consider the worst case (N=0.33); the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
Fig 8. These figures show the 1 - N AHP average case’s vulnerability score when we consider the average/expected case (N=0.165); the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
3.4 Fuzzy AHP and Fuzzy 1-N AHP
@@
Table 5: Count of Rank Reversals for Fuzzy AHP out of 100,000 simulations
@@
The simulations with rank reversals were tested to determine their potential impact on vulnerability scores. This was done by running AHP on a sample of 500 from the 4,433 cases where a rank reversal occurred, and averaging all the output layers to return the Mean Fuzzy AHP layer (Fig 9). Of these, areas that experienced changes in vulnerability due to rank reversals had high Drainage Density, or were areas that tended to receive large amounts of precipitation.
Fig 9. These figures show the Mean Fuzzy AHP’s vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
@@
For both the cases, we saw the distribution skewed to the left as compared to the original AHP (Fig 10 and Fig 11), and the vulnerability increased, which was more prominent for Case 2 (Fig 11). Due to the high weightage of the parameter, there was a increase in vulnerability scores in the areas with high population density.
Fig 10. These figures show the Fuzzy AHP Case 1 vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
Fig 11. These figures show the Fuzzy AHP Case 2’s vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
@@
Table 6: Count of First-order Rank Reversals for Fuzzy 1-N AHP out of 100,000 simulations
Reversals between Parameters Range of Original AHP results Count of Rank Reversals
@@
Table 7: Count of Second-order Rank Reversals for Fuzzy 1-N AHP out of 100,000 simulations
Reversals between Parameters Range of Original AHP results Count of Rank Reversals
@@
3.5 Comparison of analysis scores
@@
To visualize and compare the difference of the AHP layer with other approaches, we created 3 different map types (Fig 12). We created Map Type 1 by stretching the colors between the overall min max of all the different approaches, to compare the approaches among themselves. For Map Type 2 we stretched the colors between the local minimum and maximum value of each approach, to make the outliers for each individual map stand out. And for Map Type 3, we stretched the colors between the local 5 percentile and 95 percentile for each approach, to prevent the outliers from washing-out the differences in the map.
Fig 12. These figures show the comparison charts for the different variants of AHP with the original AHP across the Ganga buffer. We used three map types to highlight the differences between the approaches. Column (a) is Map type 1 (colors spread between overall min max), Column (b) is Map type 2 (colors spread between local min max), and Column (c) is Map type 3 (colors spread between 5-95 percentile). Rows 1-8 are Nested AHP, ANP, 1-N AHP (Worst case), 1-N AHP (Avg case), Fuzzy AHP (Mean), Fuzzy AHP (Case1), Fuzzy AHP (Case2) and Population density, respectively.
3.5.1 Difference between Nested AHP and ANP